Circular No. 5692

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992bm IN ANONYMOUS GALAXY M. Turatto, E. Cappellaro, and S. Benetti communicate: “We have

obtained a noisy spectrum of SN 1992bm (cf. IAUC 5672) on Jan. 15 with the Boller & Chivens spectrograph (+ CCD) at the 1.82-m reflector of the Asiago Observatory. The only clear feature is a broad (FWHM ∼ 8.0 nm) emission at 691 nm that we identify with Hα (z ∼ 0.05). This suggests that the object is probably a type-II supernova ∼ 2 months after maximum.”

GRO J0422+32 C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence,

write: “New CCD photometry of the optical counterpart of this x-ray transient has been obtained with the Haute-Provence 1.2-m telescope on Jan. 3/4, 4/5, 14/15, 15/16, and 18/19. Analysis of these data shows a very clear modulation with a period of 5.098 ± 0.005 hr and a full amplitude of 0.15 magnitude (in V). The modulation may have twice this period if the light curve is double-waved. This new period differs significantly from that reported by Kato et al. (IAUC 5676), which is incompatible with our data. The object is now 0.8 mag fainter (in V) than in early 1992 September, and substantially brighter than expected from the decline rate of 0.01 mag/day derived in the first months following the outburst (IAUC 5644, 5650). This means that the optical decay curve probably also displayed a secondary maximum in 1992 December as reported in hard x-rays by the BATSE team (IAUC 5685). Observations are continuing at Haute-Provence.”

PERIODIC COMET SCHWASSMANN-WACHMANN 1 J. Luu, University of California at Berkeley; and D. Jewitt, University

of Hawaii (UH), report: “High-resolution images reveal cyclic photometric variations that are confined to the central 1′′ of the coma. The observations were taken at the UH 2.2-m telescope in subarcsecond seeing on 1991 Nov. 17 and 1992 Nov. 27 UT. The best-fit lightcurve periods are near 10 ± 1 hr, while the range of variation is ∼ 0.5 mag. The absence of lagged variations in the near-nucleus coma suggests that the central variations are due to nucleus rotation at the stated period, rather than to coma activity. The period is short compared to the ∼ 5-day period previously estimated from coma morphological studies, but is similar to the rotation periods measured by the same technique in other short-period comets.”

1993 January 19 Daniel W. E. Green